The number of multiple births in the United States for a recent year indicated that there were 128,665 sets of twins, 7110 sets of triplets, 468 sets of quadruplets, and 85 sets of quintuplets. Choose one set of siblings at random. Find the probability that it
a. represented more than two babies
b. represented quads or quints
c. now choose now baby from these multiple births. what is the probability that the baby was a triplet?

Answer by edjones(8007)

280,957 babies
136,328 sets of siblings
.
a) (136328-128665)/136328=.0562
.
b) (468+85)/136328=.0041
.
c) (7110*3)/280957=.0759
.
Ed